# Transient and steady-state chaos in dissipative quantum systems

### v1

@@ Abstract
To probe the resulting dynamics of the open system ($\gamma=1$), we initialize it in a pure product state, $\hat{\rho}(0)$, built with the eigenstates of ^$\hat{\mathcal{H}}_{\rm TD}$, and evolve it under the Liouvillian in Eq. (2). In the Ginibre regime ($\mu=1$), the dashed line in Fig. 4(a) shows that the EE, averaged over an ensemble of random matrices, grows rapidly and saturates near the maximum value $S^{\rm max}_{\rm en}=\frac{1}{2}\ln(N)$, indicating both short-time and steady-state chaos. In contrast, for the 2d Poisson statistics ($\mu=0$), the EE grows slowly and saturates at a much lower value, consistent with the absence of chaos. The steady-state density matrix in this regime retains a significant degree of purity and can be expressed as $\hat{\rho}_{\rm R}=\sum_{i}\eta_{i}|{u_{i}}\rangle\langle{u_{i}}|$, where only a few eigenvalues $\eta_{i}$ are significantly large.
@@
To decouple transient chaos from steady-state behavior, we consider a second choice of ^$\hat{\mathcal{H}}_{\rm I}$ by deforming the chaotic Hamiltonian via ^$\hat{\mathcal{H}}_{\rm I}=\hat{\mathcal{P}}(\chi)\hat{\mathcal{H}}_{\rm GOE}\hat{\mathcal{P}}(\chi)$, where $\hat{\mathcal{P}}(\chi)=\mathbb{I}-\chi(|{u_{1}}\rangle\langle{u_{1}}|+|{u_{2}}\rangle\langle{u_{2}}|)$ is a projection operator that protects the dominant eigenstates $|{u_{1}}\rangle$ and $|{u_{2}}\rangle$ of the regular steady-state density matrix $\hat{\rho}_{\rm R}$ from mixing with the random part. The parameter $\chi\in[0,1]$ controls the strength of this suppression and allows for retention of partial purity in the steady state.
@@
Figure 4: Transient and steady-state chaos for the random matrix toy model. (a) Time evolution of the entanglement entropy $\mathcal{S}_{\rm en}$, averaged over an ensemble of random matrices, for the Liouvillian in Eq. (2) with $\mu=0$ (solid red line), resulting in 2d Poisson level statistics, with $\mu=1$, $\chi=0$ (dashed line), resulting in Ginibre spectral statistics, and the projected case with $\mu=1$, $\chi=1$ (solid blue line), also resulting in Ginibre spectral statistics. (b) Evolution of $\mathcal{S}_{\rm en}$ for different values of the deformation parameter $\chi$ and fixed $\mu=1$. The horizontal dashed line in (a)-(b) indicates the maximal entanglement entropy, $S^{\rm max}_{\rm en}=\frac{1}{2}\ln(N)$. We use $N=49$ and $\gamma=1$. The random numbers are drawn from a Gaussian distribution with zero mean and unit standard deviation.
@@
This deformation preserves the Ginibre spectral statistics at $\mu=1$ for all values of $\chi$ (see SM [54]), and it also leads to the rapid initial growth of the EE, as shown by the blue solid line in Fig. 4(a). However, the long-time entropy saturates well below the maximum value $S^{\rm max}_{\rm en}=\frac{1}{2}\ln(N)$, indicating the absence of chaotic behavior in the steady state. This behavior is consistent with the scenario of transient chaos.
@@
Figure 4(b) further demonstrates that while the short-time growth of entropy is largely unaffected by the projector, the long-time saturation value decreases monotonically with increasing $\chi$. This behavior closely mirrors that of the open ADM, where dissipation suppresses long-time chaos but preserves transient chaos. These results show that Ginibre spectral statistics is directly associated with transient chaos, but does not necessarily reflect steady-state chaos. This resolves the perceived breakdown of the quantum-classical correspondence reported in [26].
@@
Conclusions.– This work demonstrated that chaos in open quantum systems manifests differently across timescales and cannot be fully characterized by spectral statistics alone. Using the open anisotropic Dicke model, we showed that entanglement entropy and OTOCs reliably distinguish between transient and steady-state chaos, restoring the quantum-classical correspondence in the presence of dissipation at both short and long times. In contrast, Ginibre spectral statistics of the Liouvillian spectrum are connected to early-time chaotic behavior and therefore cannot differentiate between the two asymptotic regimes, as supported by our random matrix toy model.
@@
Acknowledgements.– L. F. S. thanks David Villaseñor and Pablo Barberis-Blostein for useful discussions and support from the National Science Foundation Center for Quantum Dynamics on Modular Quantum Devices (CQD-MQD) under Award Number 2124511.
@@
This supplemental material provides additional figures and analyses that support the discussions in the main text. In Sec. I, we briefly examine the classical dynamics of the open anisotropic Dicke model (ADM), focusing on the phase diagram of non-equilibrium phases and the emergence of dissipative chaos. In Sec. II, steady-state chaos is analyzed with the saturation value of the entanglement entropy (EE) and standard deviations of the $z$ component of the spin operator and photon number. Finally, Sec. III presents results for the spectral statistics of the Liouvillian associated with the random matrix toy model.
@@ I Classical dynamics in the open anisotropic Dicke model
$\frac{d\langle\hat{\mathcal{O}}\rangle}{dt}={\rm Tr}(\hat{\mathcal{O}}\dot{\hat{\rho}}).$
In the limit $S\rightarrow\infty$, the scaled operators $\hat{a}/\sqrt{S}=(\hat{x}+i\hat{p})/\sqrt{2}$ and $\hat{\vec{s}}=\hat{\vec{S}}/S$ behave classically, since they satisfy the commutation relations $[\hat{x},\hat{p}]=i/S$ and $[\hat{s}_{a},\hat{s}_{b}]=i\epsilon_{abc}\hat{s}_{c}/S$, where $1/S$ plays the role of a reduced Planck constant. The classical equations of motion for the scaled observables can be derived from the master equation and are given by
@@
To understand the overall dynamics and describe the different non-equilibrium phases, we analyze the fixed points and attractors of the equations of motion in Eq. (S3). The various phases of the above model for the coupling parameters $\lambda_{-}$ and $\lambda_{+}$ and a fixed dissipation strength $\kappa$ are summarized in the phase diagram, in Fig. S5(a).
Figure S5: The classical dynamics of the open anisotropic Dicke model. (a) The classical phase-diagram on the $\lambda_{-}-\lambda_{+}$. (b) The spin dynamics over the Bloch sphere at different values of $\lambda_{+}$ for $\lambda_{-}=2.0$. All energies (time) are measured by $\omega(1/\omega)$. We set $\omega=1.0,\omega_{0}=1.0$ and dissipation strength $\kappa=1$.
For the ADM, there are two types of normal phases, NP and NP, characterized by vanishing photon number $n^{*}=0$ and spin polarization $s_{z}^{*}=-1$ and $s_{z}^{*}=+1$, respectively. Fixing $\lambda_{-}=2$ and following the vertical line in Fig. S5(a) by increasing $\lambda_{+}$, the normal phase NP undergoes a transition to a superradiant phase (SR) with non-zero photon number $n^{*}\neq 0$ and spin polarization $|s_{z}^{*}|<1$. Then the SR phase becomes unstable at a critical coupling and undergoes a Hopf bifurcation, giving rise to a limit cycle (LC). Further increasing $\lambda_{+}$ leads to the instability of the limit cycle and emergence of chaotic motion, which can be quantified with the Lyapunov exponent [60, 61, 62]. There is a narrow region above the chaotic domain in the phase-diagram, where the SR phase regains its stability and coexists with the stable normal phase NP. As the interaction gets even stronger, the SR phase disappears entirely, and only the normal phase NP remains as stable steady state. The detailed description of these non-equilibrium phases and the dissipative dynamics of the open ADM is presented in Ref. [37]. The spin dynamics across distinct dynamical regimes for different values of $\lambda_{+}$ with fixed $\lambda_{-}$ are illustrated on the Bloch sphere in Fig. S5(b).
@@ II Steady-state chaos in the open anisotropic Dicke model
As shown in Fig. S5(a), the open ADM exhibits chaos within a narrow triangular region of the phase-diagram, where no stable attractor persists. The presence of chaos can be quantified using the saturation value of the classical Lyapunov exponent $\Lambda_{\rm ss}$, as shown in Fig. 2(c) of the main text. Quantum mechanically, such steady-state chaos can be captured by the saturation value of the total entanglement entropy, $\mathcal{S}_{\rm ss}^{\rm en}$, and the standard deviations of physical quantities, such as the $z$ component of the spin operator, $(\Delta S_{z})_{\rm ss}$, and the photon number, $(\Delta n)_{\rm ss}$. For a fixed coupling strength $\lambda_{-}=2$, we examine in Fig. S6 how the values of the three quantities change as $\lambda_{+}$, corresponding to the vertical line in Fig. S5(a). We see that all three quantities reach their maximum values in the chaotic regime.
Figure S6: Signature of steady-state chaos based on the saturation value of the (a) total entanglement entropy $\mathcal{S}_{\rm ss}^{\rm en}$, (b) fluctuation of the $z$ component of spin operator $(\Delta S_{z})_{\rm ss}$, and (c) fluctuation of the photon number $(\Delta n)_{\rm ss}$ as a function of $\lambda_{+}$ for $\lambda_{-}$ = 2. We consider the parameters: $\omega=\omega_{0}=1,\kappa=1,S=5$.
@@ III Toy model using random matrices
In the main text, we explained that the spectral statistics of the Liouvillian spectrum cannot distinguish transient from steady-state chaos. Here, we present the results for the Liouvillian spectrum of the random matrix toy model introduced in the main text.
The total Hamiltonian is defined as
^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\rm TD}+\frac{\mu}{\sqrt{N}}\hat{\mathcal{H}}_{\rm I},$
where ^$\hat{\mathcal{H}}_{\rm TD}$ is an $N\times N$ tridiagonal random matrix representing a regular system with Poisson spectral statistics, and ^$\hat{\mathcal{H}}_{\rm I}$ introduces level repulsion, controlled by the parameter $\mu$. Dissipation is incorporated with a jump operator $\sqrt{\gamma}\hat{L}$ where $\hat{L}_{m+1,m}$ are nonzero real random elements for $m=1,2,3...$ and $\gamma$ sets the strength of dissipation. The corresponding Liouvillian superoperator is given in Eq. (2) of the main text. The perturbation Hamiltonian ^$\hat{\mathcal{H}}_{\rm I}$ is constructed as
^$\hat{\mathcal{H}}_{\rm I}(\chi)$ $=$ $\hat{\mathcal{P}}(\chi)\hat{\mathcal{H}}_{\rm GOE}\hat{\mathcal{P}}(\chi)$ (S4)
@@
where ^$\hat{\mathcal{H}}_{\rm GOE}$ is a random matrix from a Gaussian orthogonal ensemble (GOE) and $\hat{\mathcal{P}}(\chi)$ is the projection operator that suppresses dominant eigenstates $|{u_{1}}\rangle$ and $|{u_{2}}\rangle$ of the steady-state density matrix $\hat{\rho}_{\rm R}$ of the regular Liouvillian with $\mu=0$ and same jump operator. The parameter $\chi\in[0,1]$ controls the strength of the projection. When $\chi=0$, there is no projection and ^$\hat{\mathcal{H}}_{\rm GOE}$.
@@
At $\mu=1$, the spectrum displays Ginibre statistics regardless of the value of $10<\chi<$ 1, as seen for $\chi=0$ in Fig. S7(b) and $\chi=1$ in Fig. S7(c). This contrasts with the steady state associated with the zero eigenvalue of the Liouvillian, which shows strong dependence on $\chi$, as shown in Fig. 4(b) of the main text. For large $\chi$, the system retains substantial purity at long times, and the EE saturates below its maximal value, indicating suppression of steady-state chaos despite early-time signatures of chaos. These results confirm that Ginibre spectral correlations in the Liouvillian spectrum are directly connected to rapid short-time entanglement growth, independent of the steady-state behavior, highlighting the limitations of spectral statistics in fully diagnosing chaos in open quantum systems.

### v2

@@ Abstract
In the Ginibre regime ($\mu=1$), the dashed line in Fig. 4(a) shows that the VNE of the reduced density matrix of a subsystem $\mathcal{S}_{\rm VN}$, averaged over an ensemble of random matrices, grows rapidly and saturates near the maximum value $\mathcal{S}^{\rm max}_{\rm VN}=\frac{1}{2}\ln(N)$, indicating both short-time and steady-state chaos. In contrast, for the 2d Poisson statistics ($\mu=0$), the entropy grows slowly and saturates at a much lower value, consistent with the absence of chaos. The steady-state density matrix in this regime retains a significant degree of purity and can be expressed as $\hat{\rho}_{\rm R}=\sum_{i}\eta_{i}|{u_{i}}\rangle\langle{u_{i}}|$, where only a few eigenvalues $\eta_{i}$ are significantly large.
@@
To decouple transient chaos from steady-state behavior, we consider a second choice of $\hat{\mathcal{H}}_{\rm I}$ by deforming the chaotic Hamiltonian via $\hat{\mathcal{H}}_{\rm I}=\hat{\mathcal{P}}(\chi)\hat{\mathcal{H}}_{\rm GOE}\hat{\mathcal{P}}(\chi)$, where $\hat{\mathcal{P}}(\chi)=\mathbb{I}-\chi(|{u_{1}}\rangle\langle{u_{1}}|+|{u_{2}}\rangle\langle{u_{2}}|)$ is a projection operator that protects the dominant eigenstates $|{u_{1}}\rangle$ and $|{u_{2}}\rangle$ of the regular steady-state density matrix $\hat{\rho}_{\rm R}$ from mixing with the random part. The parameter $\chi\in[0,1]$ controls the strength of this suppression and allows for retention of partial purity in the steady state.
@@
Figure 4: Transient and steady-state chaos for the random matrix toy model. (a) Time evolution of the von Neumann entropy $\mathcal{S}_{\rm VN}$, averaged over an ensemble of random matrices, for the Liouvillian in Eq. (2) with $\mu=0$ (solid red line), resulting in 2d Poisson level statistics, with $\mu=1$, $\chi=0$ (dashed violet line), resulting in Ginibre spectral statistics, and the projected case with $\mu=1$, $\chi=1$ (solid blue line), also resulting in Ginibre spectral statistics. (b) Evolution of $\mathcal{S}_{\rm VN}$ for different values of the deformation parameter $\chi$ and fixed $\mu=1$. The horizontal dashed line in (a)-(b) indicates the maximal entropy of the subsystem, $\mathcal{S}^{\rm max}_{\rm VN}=\frac{1}{2}\ln(N)$. We use $N=49$ and $\gamma=1$. The random numbers are drawn from a Gaussian distribution with zero mean and unit variance.
@@
This deformation preserves the Ginibre spectral statistics at $\mu=1$ for all values of $\chi$ (see SM [55]), and it also leads to the rapid initial growth of the VNE, as shown by the blue solid line in Fig. 4(a). However, the long-time entropy saturates well below the maximum value $\mathcal{S}^{\rm max}_{\rm VN}$, indicating absence of chaotic behavior in the steady state. This behavior is consistent with the scenario of transient chaos.
@@
Figure 4(b) further demonstrates that while the short-time growth of the entropy is largely unaffected by the projector, the long-time saturation value decreases monotonically with increasing $\chi$. This behavior closely mirrors that of the open ADM, where dissipation suppresses long-time chaos but preserves transient chaos. These results show that Ginibre spectral statistics is directly associated with transient chaos, but does not necessarily reflect steady-state chaos. This resolves the perceived breakdown of the quantum-classical correspondence reported in [29].
@@
Conclusions.– This work demonstrated that chaos in open quantum systems manifests differently across timescales and cannot be fully characterized by spectral statistics alone. Using the open anisotropic Dicke model, we showed that the von Neumann entropy of subsystems of the model and OTOCs reliably distinguish between transient and steady-state chaos, restoring the quantum-classical correspondence in the presence of dissipation at both short and long times. In contrast, Ginibre spectral statistics of the Liouvillian spectrum are connected to early-time chaotic behavior and therefore cannot differentiate between the two asymptotic regimes, as supported by our random matrix toy model.
@@
Acknowledgments.– L. F. S. thanks David Villaseñor and Pablo Barberis-Blostein for useful discussions and support from the National Science Foundation Center for Quantum Dynamics on Modular Quantum Devices (CQD-MQD) under Award Number 2124511.
@@
This supplemental material provides additional figures and analyses that support the discussions in the main text. In Sec. I, we briefly examine the classical dynamics of the open anisotropic Dicke model (ADM), focusing on the phase diagram of non-equilibrium phases and the emergence of dissipative chaos. In Sec. II, steady-state chaos is analyzed with the saturation value of the von Neumann entropy (VNE) and standard deviations of the $z$ component of the spin operator and the photon number. Analysis of the dependence on the spin size is also provided. Finally, Sec. III presents results for the spectral statistics of the Liouvillian associated with the random matrix toy model.
@@ I Classical dynamics in the open anisotropic Dicke model
Figure S5: The classical dynamics of the open anisotropic Dicke model. (a) The classical phase-diagram on the $\lambda_{-}-\lambda_{+}$ plane. (b) The spin dynamics over the Bloch sphere at different values of $\lambda_{+}$ for $\lambda_{-}=2.0$. All energies (time) are measured by $\omega(1/\omega)$. We set $\omega=1.0,\omega_{0}=1.0$ and dissipation strength $\kappa=1$.
@@
[…]
In the limit $S\rightarrow\infty$, the scaled operators $\hat{a}/\sqrt{S}=(\hat{x}+i\hat{p})/\sqrt{2}$ and $\hat{\vec{s}}=\hat{\vec{S}}/S$ behave classically, since they satisfy the commutation relations $\hat{x},\hat{p}]=i/S$ and $\hat{s}_{a},\hat{s}_{b}]=i\epsilon_{abc}\hat{s}_{c}/S$, where $1/S$ plays the role of a reduced Planck constant. The classical equations of motion for the scaled observables can be derived from the master equation and are given by
@@
To understand the overall dynamics and describe the different non-equilibrium phases, we analyze the fixed points and attractors of the equations of motion in Eqs. (S3). The various phases of the above model for the coupling parameters $\lambda_{-}$ and $\lambda_{+}$ and a fixed dissipation strength $\kappa$ are summarized in the phase diagram, in Fig. S5(a).
For the ADM, there are two types of normal phases, NP and NP, characterized by vanishing photon number $n^{*}=0$ and spin polarization $s_{z}^{*}=-1$ and $s_{z}^{*}=+1$, respectively. Fixing $\lambda_{-}=2$ and following the vertical line in Fig. S5(a) by increasing $\lambda_{+}$, the normal phase NP undergoes a transition to a superradiant phase (SR) with non-zero photon number $n^{*}\neq 0$ and spin polarization $|s_{z}^{*}|<1$. Then the SR phase becomes unstable at a critical coupling and undergoes a Hopf bifurcation, giving rise to a limit cycle (LC). Further increasing $\lambda_{+}$ leads to the instability of the limit cycle and emergence of chaotic motion, which can be quantified with the Lyapunov exponent [65, 66, 67]. There is a narrow region above the chaotic domain in the phase-diagram, where the SR phase regains its stability and coexists with the stable normal phase NP. As the interaction gets even stronger, the SR phase disappears entirely, and only the normal phase NP remains as stable steady state. The detailed description of these non-equilibrium phases and the dissipative dynamics of the open classical ADM is presented in Ref. [41]. The spin dynamics across distinct dynamical regimes for different values of $\lambda_{+}$ with fixed $\lambda_{-}$ are illustrated on the Bloch sphere in Fig. S5(b).
@@ II Steady-state chaos in the open anisotropic Dicke model
As shown in Fig. S5(a), the open ADM exhibits chaos within a narrow triangular region of the phase-diagram, where no stable attractor persists. The presence of chaos can be quantified using the saturation value of the classical Lyapunov exponent $\Lambda_{\rm ss}$, as shown in Fig. 2(c) of the main text. Quantum mechanically, such steady-state chaos can be captured by the saturation value of the total VNE, $\mathcal{S}_{\rm ss}^{\rm VN}$, and the standard deviations of physical quantities, such as the $z$ component of the spin operator, $(\Delta S_{z})_{\rm ss}$, and the photon number, $(\Delta n)_{\rm ss}$, extracted from the corresponding FOTOCs. For a fixed coupling strength $\lambda_{-}=2$, we examine in Fig. S6 how the values of these three quantities change as $\lambda_{+}$ increases, corresponding to the vertical line in Fig. S5(a). We see that all three quantities reach their maximum values in the chaotic regime.
Figure S6: Signature of steady-state chaos based on the saturation value of the (a) total von Neumann entropy $\mathcal{S}_{\rm ss}^{\rm VN}$, (b) fluctuation of the $z$ component of the spin operator $(\Delta S_{z})_{\rm ss}$, and (c) fluctuation of the photon number $(\Delta n)_{\rm ss}$ as a function of $\lambda_{+}$ for $\lambda_{-}$ = 2. We consider the parameters: $\omega=\omega_{0}=1,\kappa=1,S=5$.
II.1 Dependence on the spin size
As discussed in the main text, steady-state chaos in the open quantum ADM is effectively captured by the relative enhancement of the subsystem von Neumann entropy, $\mathcal{S}^{\rm VN}_{\rm ss}$, as illustrated in the color-scale phase diagram of Fig.2 (d) of the main text. Here, we focus on the suppression of chaos associated with the emergence of the stable normal phase NP as the coupling strength $\lambda_{+}$ increases. Figure S7 shows that $\mathcal{S}^{\rm VN}_{\rm ss}$ decreases gradually with increasing $\lambda_{+}$, signaling a crossover from the chaotic to the regular regime. The relative difference in VNE between these regimes grows with the spin magnitude $S$, yielding a sharper quantum distinction between chaotic and regular dynamics.
To quantify this contrast, we define
$\Delta\mathcal{S}^{\rm VN}_{\rm ss}=\mathcal{S}^{\rm VN}_{\rm ss}(\lambda_{+}^{(1)})-\mathcal{S}^{\rm VN}_{\rm ss}(\lambda_{+}^{(2)}),$
where the coupling strengths $\lambda_{+}^{(1)}$ and $\lambda_{+}^{(2)}$ lie in the chaotic and regular regions, respectively (marked by the vertical dashed lines in Fig. S7). As shown in the inset of this figure, $\Delta\mathcal{S}^{\rm VN}_{\rm ss}$, making the separation between chaotic and regular regimes increasingly pronounced with $S$.
Figure S7: Crossover from the chaotic to the regular regime. The main panel shows how the steady state value of total VNE, $\mathcal{S}^{\rm VN}_{\rm ss}$, changes with the coupling $\lambda_{+}$ for spin magnitudes $S=5$ and $10$. The red shaded area indicates the chaotic regime. The inset shows the difference $\Delta\mathcal{S}^{\rm VN}_{\rm ss}$ in the values of the total VNE at the couplings $\lambda_{+}^{(1)}$ (chaotic region) and $\lambda_{+}^{(2)}$ (regular region) as a function of spin magnitude $S$. Chosen parameters: $\omega=\omega_{0}=1,\lambda_{-}=2,\kappa=1$.
@@ III Toy model using random matrices
In the main text, we explained that the spectral statistics of the Liouvillian spectrum cannot distinguish transient from steady-state chaos. Here, we present the results for the Liouvillian spectrum of the random matrix toy model introduced in the main text. The total Hamiltonian is defined as
$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\rm TD}+\frac{\mu}{\sqrt{N}}\hat{\mathcal{H}}_{\rm I},$
where $\hat{\mathcal{H}}_{\rm TD}$ is an $N\times N$ tridiagonal random matrix representing a regular system with Poisson spectral statistics, and $\hat{\mathcal{H}}_{\rm I}$ introduces level repulsion, controlled by the parameter $\mu$. Dissipation is incorporated with a jump operator $\sqrt{\gamma}\hat{L}$ where $\hat{L}_{m+1,m}$ are nonzero real random elements for $m=1,2,3$... and $\gamma$ sets the strength of dissipation. The corresponding Liouvillian superoperator is given in Eq. (2) of the main text. The perturbation Hamiltonian $\hat{\mathcal{H}}_{\rm I}$ is constructed as
$\hat{\mathcal{H}}_{\rm I}(\chi)$ $=$ $\hat{\mathcal{P}}(\chi)\hat{\mathcal{H}}_{\rm GOE}\hat{\mathcal{P}}(\chi)$ (S4)
@@
where $\hat{\mathcal{H}}_{\rm GOE}$ is a random matrix from a Gaussian orthogonal ensemble (GOE) and $\hat{\mathcal{P}}(\chi)$ is the projection operator that suppresses dominant eigenstates $|{u_{1}}\rangle$ and $|{u_{2}}\rangle$ of the steady-state density matrix $\hat{\rho}_{\rm R}$ of the regular Liouvillian with $\mu=0$ and same jump operator. The parameter $\chi\in[0,1]$ controls the strength of the projection. When $\chi=0$, there is no projection and $\hat{\mathcal{H}}_{\rm I}=\hat{\mathcal{H}}_{\rm GOE}$.
@@
At $\mu=1$, the spectrum displays Ginibre statistics regardless of the value of $0<\chi<1$, as seen for $\chi=0$ in Fig. S8(b) and $\chi=1$ in Fig. S8(c). This contrasts with the steady state associated with the zero eigenvalue of the Liouvillian, which shows strong dependence on $\chi$, as shown in Fig. 4(b) of the main text. For large $\chi$, the system retains substantial purity at long times, and the VNE saturates below its maximal value, indicating suppression of steady-state chaos despite early-time signatures of chaos. These results confirm that Ginibre spectral correlations in the Liouvillian spectrum are directly connected to rapid short-time entropy growth, independent of the steady-state behavior, highlighting the limitations of spectral statistics in fully diagnosing chaos in open quantum systems.